OpenStudy (anonymous):

When the equation 3x - 4y = 24 is written in slope-intercept form, what is the value of m?

OpenStudy (anonymous):

Transform the equation so it looks like Y = mX + b

OpenStudy (anonymous):

3x-4y=24 Since 3x does not contain the variable to solve for, move it to the right-hand side of the equation by subtracting 3x from both sides. -4y=-3x+24 Divide each term in the equation by -4. -(4y)/(-4)=-(3x)/(-4)+(24)/(-4) Simplify the left-hand side of the equation by canceling the common factors. y=-(3x)/(-4)+(24)/(-4) Simplify the right-hand side of the equation by simplifying each term. y=(3x)/(4)-6 To find the slope and y intercept, use the y=mx+b formula where m=slope and b is the y intercept. y=mx+b Using the y=mx+b formula, m=(3)/(4) and b=-6. m=(3)/(4), b=-6

OpenStudy (anonymous):

Thank you so much!!

OpenStudy (anonymous):

Medal? :)

OpenStudy (anonymous):

your welcome buddy